### v1

@@ Abstract
Context. The search of exoplanets around nearby M dwarfs is a crucial milestone to perform the census of planetary systems in the vicinity of our Solar System.
Aims. Since 2018 our team is carrying a blind search program for planets around nearby M dwarfs with the near-IR spectro-polarimeter and velocimeter SPIRou at the CFHT and the optical velocimeter SOPHIE at the OHP in France. The aim of this paper is to present our results on Gl 410, a 0.55 M⊙ 480150 Myr old active M dwarf distant 12 pc.
Methods. We search for planetary companions using radial velocities (RVs). We use the line-by-line (LBL) technique to measure the RVs with SPIRou and the template matching method with SOPHIE. Three different methods, two based in principal component analysis (PCA), are used to clean the SPIRou RVs for systematics. Gaussian processes (GP) modeling is applied to correct the SOPHIE RVs for stellar activity. The and apodize sine periodogram (ASP) analysis is used to search for planetary signals in the SPIRou data taking into account activity indicators. We analyse TESS data and search for planetary transits.
Results. We report the detection of a sin=8.41.3 M⊕ sub-Neptune planet at a period of 6.0200.004 days in circular orbit with SPIRou. The same signal, although with lower significance, is also retrieved in the SOPHIE RV data after correction for activity using a GP trained on SPIRou’s longitudinal magnetic field () measurements. TESS data indicates that the planet is not transiting. We find within the SPIRou wPCA RVs tentative evidence for two additional planetary signals at 2.99 and 18.7 days.
Conclusions. Infrared RVs are a powerful method to detect extrasolar planets around active M dwarfs. Care should be taken however to correct/filter systematics generated by residuals of the telluric correction or small structures in the detector plane. The LBL technique combined with PCA offers a promising way to reach this objective. Further monitoring of Gl 410 is necessary.
@@
One of the most interesting open questions in current Astronomy is the census of extrasolar planets in the Solar System vicinity. The identification of closest extrasolar planetary systems is a crucial milestone for the study of exoplanet atmospheres and biomarkers using future instrumentation combining high-dispersion spectroscopy with high contrast imaging (e.g., Kasper et al., 2021; Chauvin, 2024). As the most numerous stars in the Galaxy, thus in Solar neighbourhood, are low-mass stars ( M⊙), in recent years, there has been a growing interest in extending the planet search effort from Solar-like stars to M dwarfs.
@@
Radial velocity (Bonfils et al., 2013; Pinamonti et al., 2022, Mignon et al. 2024 sub.) and transit surveys (Dressing & Charbonneau, 2013; Mulders et al., 2015; Dressing & Charbonneau, 2015; Hsu et al., 2020; Ment & Charbonneau, 2023) have revealed that a large fraction of M dwarfs, if not all, host a planetary system, and have provided strong evidence supporting the hypothesis that the rocky planet ratio per star is indeed larger than one. Extrapolating the results from field stars to our stellar vicinity is extremely encouraging, as it suggests that is highly likely that a large fraction of nearby M dwarfs host multi-planetary systems. Nearby stars are distributed all over the sky at random orientations. A small fraction of sources are expected to present planets in transit. Missions such as TESS and in the future PLATO will detect a significant fraction of those transiting planets. Nevertheless, to establish the planetary systems census of our closest stellar neighbours, techniques such as astrometry or radial velocities (RV) need to be used in addition to transits.
@@
Surveys for planetary systems around nearby bright M dwarfs with high-precision velocimeters have been carried so far mainly in the optical domain. Recently the RV search has been extended to the near infrared (IR) domain, with the start of operations of new velocimeters such as HPF (Mahadevan et al., 2014), CARMENES (Quirrenbach et al., 2018), SPIRou (Donati et al., 2020) and NIRPS (Bouchy et al., 2017). High-precision IR RVs enables us to explore lower mass stars, sometimes inaccessible for precise RV work in the optical due to their faintness. Furthermore, IR RVs have the potential of being less affected by stellar activity than optical RVs specifically for highly or moderately active stars (e.g. Carmona et al., 2023).
@@
With more than two-thousand low-mass M dwarfs in the 25 pc vicinity, the effort to monitor all these objects is colossal. Thanks to large programmes with the near-IR spectra-polarimeter and velocimeter SPIRou at CHFT (the SPIRou Legacy Survey, the SPICE large programme) and a series of PI SPIRou follow-up programs our team has been monitoring a selected sub-sample of around 60 nearby M dwarfs. At the time of writing, we have obtained more than 100 visits per source for around 45 of them.
@@
In this paper, we report the detection of a sub-Neptune mass planet in a 6 day period around the nearby (12 pc) 0.55 M⊙ 480 Myr old M dwarf Gl 410 with SPIRou, and the tentative detection of two additional planetary signals at 2.99 and 18.7 days. We complement our near-IR data set with quasi simultaneous optical observations obtained with the velocimeter SOPHIE. The paper is organised as follows. We start describing the physical properties of Gl 410 in Sect. 2; then, in Sect. 3, we present the details of the optical and infrared RV measurements. In Sect. 4, we analyze our RV measurements and test the robustness of the planet detection. Then in Sect. 5, we search for the 6 day period signal detected with SPIRou in the SOPHIE data using Gaussian Processes activity filtering techniques. We derive the orbital and physical properties of the detected 6.02 day planet using MCMC techniques in Sect. 6. In Sect. 7, we investigate the presence of other planetary signals within the SPIRou wPCA RVs. In Sect. 8, we present TESS observations of Gl 410 and investigate whether Gl 410b is transiting. We discuss our findings in Sect. 9 and finally, in Sect. 10, we provide our conclusions.
2 Gl410
We summarize the basic physical properties of Gl 410 (DS Leo, HD 95650) in Table 1. Gl 410 has a spectral type M1.0V and a mass of 0.550.02 M⊙ (Cristofari et al., 2022). It is a nearby star located at a distance of 11.9327 pc (Gaia DR3, Bailer-Jones et al., 2021; Gaia Collaboration et al., 2023). Gl 410 is known to exhibit cyclic photometric variability in the optical with periods of about 14 d which is interpreted as caused by rotational modulation of starspots (Fekel & Henry, 2000). The rotational period of Gl 410 has been tightly constrained using spectropolarimetry observations in the optical with ESPaDOnS, NARVAL and HARPS (Donati et al., 2008; Hébrard et al., 2016), and the near-IR with SPIRou (Fouqué et al., 2023; Donati et al., 2023). The latest estimation of is of 13.910.09 days (Donati et al., 2023). Gl 410 displays surface differential rotation with periods at the equator and pole of 13.370.86 and 14.961.25 days respectively (Hébrard et al., 2016). Based on gyrochronogly Fouqué et al. (2023) derived an age of 0.890.1 Gyr for Gl 410. However, there is the possibility that Gl 410 is in fact younger. In Sect. A in the Appendix, we provide further details. We find that empirical rotation-temperature sequences in clusters of known ages indicate that Gl 410 is 480150 Myr, which ranks it among the youngest stars in the solar neighbourhood (Silva Aguirre et al., 2018; Chen et al., 2020), and suggest that Gl 410 might be related to the Ursa Major (UMA) cluster which is approximately 400 Myr old.
Cristofari et al. (2023) based on SPIRou observations measured a average magnetic small scale field strength ¡¿ of 0.71 0.03 kG for Gl 410. Together with the M dwarfs GJ 1289 and GJ 1286, Gl 410 is part of most magnetic stars among the M dwarfs studied in the ¡¿ surveys of Reiners et al. (2022) and Cristofari et al. (2023). The magnetic activity of Gl 410 is further indicated by the presence of emission in the optical spectra. Gl 410 is part of the top one third of the sample in terms of emission strength among the 331 M-stars surveyed by Schöfer et al. (2019).
Zeeman Doppler Imaging (ZDI) investigations by Donati et al. (2008), Hébrard et al. (2016) and Bellotti et al. (2024) show that the large-scale magnetic field of Gl 410 rapidly evolves both in strength and in topology. Between 2007 and 2008, the field had a predominantly toroidal geometry with an average strength of 100 G (Donati et al., 2008). In 2014, the field remained mostly-toroidal with axisymmetric geometry but the average field strength decreased to 60 G (Hébrard et al., 2016). In 2020b to 2021a the magnetic energy became almost equally distributed between the poloidal and the toroidal components and the average field strength decreased to 44 G (Bellotti et al., 2024). In the latest measurements from 2021b to 2022a, the poloidal component became much stronger than the toroidal component with 73% of the magnetic energy and the average field strength further decreased to 18 G (Bellotti et al., 2024).
@@ 3.1 SOPHIE
Gl 410 was observed during two periods 9 years apart with the high-resolution optical velocimeter SOPHIE222Spectrographe pour l’Observation des Phénomènes des Intérieurs stellaires et des Exoplanètes, meaning, spectrograph for the observation of the phenomena of the stellar interiors and of the exoplanets. (Perruchot et al., 2008; Bouchy et al., 2013) at the 1.93m Telescope of the Haute Provence Observatory in the South of France. The first period with 45 visits was observed from March 14, 2010, to June 14, 2012. A second observation period was planned to coincide with the SPIRou observations, with 58 visits performed from January 26, 2021, to March 04, 2023. Spectra were reduced using the SOPHIE’s data reduction system (Bouchy et al., 2009; Heidari, 2022). RV measurements were obtained using a template-matching technique implemented in the code NAIRA (Astudillo-Defru et al., 2015) which has been adapted for SOPHIE (Hobson et al., 2018). We provide a summary of the SOPHIE the RV measurements in the Table 10 in the Appendix. In Table 2, we provide a summary of the properties of the RV time-series. Figure 1 displays the median subtracted SOPHIE RV data.
@@ 3.2 SPIRou
Gl 410 was observed with the near-IR velocimeter and polarimeter SPIRou333SPectropolarimètre InfraRouge. (Donati et al., 2020) mounted at the Canada-France-Hawaii Telescope atop Mauna-Kea during 156 visits from April 2018 to July 2023. Data were obtained in the context of the SPIRou’s commissioning in 2018, the SPIRou Legacy Survey from 2019 to 2022A, and a series of PI follow-program programs (PI. Carmona, PI. Artigau) in 2022B, 2023A and 2023B. SPIRou nightly observations typically consist of a Stokes V sequence of four exposures, each taken with different positions of the Fresnel rhombs inside the polarimeter. The combination of the channel A and channel B of the four sub-exposures enables to recover the Stokes V (and Stokes I) spectra for the magnetic field analysis. For each science observation, a simultaneous Fabry-Pérot exposure is obtained in the calibration channel C. In the context of the radial velocity analysis, we use the combined extraction of the science channels A and B and each sub-exposure of the Stokes V sequence is treated independently.
For the radial velocity analysis we used the data reduced using the APERO (Cook et al., 2022) pipeline version v07275 installed at the SPIRou Data Centre hosted at the Laboratoire d’Astrophysique de Marseille. APERO extracts the combined science channel A and channel B spectra, performs the wavelength calibration using nightly observations of the Uranium-Neon lamp and Fabry-Pérot exposures, and do the telluric correction. The telluric correction uses the combination of an atmospheric absorption model generated by TAPAS (Bertaux et al., 2014), a empirical template of the star deduced from observations at different BERV, and a library of nightly observations of telluric standard stars taken since the commissioning of the instrument in 2018. Further details on the telluric correction are given in Artigau et al. (2014), Cook et al. (2022) and Artigau et al. (in prep.). Individual radial-velocity measurements were obtained using the LBL technique (Artigau et al., 2022) and they are corrected by the BERV and the drift of the Fabry-Pérot at each exposure. A further correction for the instrumental radial-velocity zero point is done using a Gaussian Process modeling based on observations of a sample of radial-velocity standard stars observed by SPIRou (Cadieux et al., 2022).
For the analysis of the stellar activity, we used the simultaneous longitudinal magnetic () field measurements obtained with the Libre-ESpRIT pipeline444The nominal ESPaDOnS data reduction pipeline (Donati et al., 1997) adapted for the SPIRou data (Donati et al., 2020). previously published in Donati et al. (2023).
@@ 3.3 TESS
The Transit Exoplanet Survey (TESS, Ricker et al., 2014) observed Gl410 (TIC 97488127) with two-minute cadence during Sector 22, from February 18th to March 18th of 2020. We obtained the Presearch Data Conditioning (PDC) flux time-series (Smith et al., 2012; Stumpe et al., 2012, 2014) processed by the TESS Science Processing Operations Center (SPOC, Jenkins et al., 2016) from the Mikulski Archive for Space Telescopes (MAST)555mast.stsci.edu.

### v2

@@ Abstract
Context. The search for exoplanets around nearby M dwarfs represents a crucial milestone in the census of planetary systems in the vicinity of our Solar System.
Aims. Since 2018 our team has been conducting a blind search program for planets around nearby M dwarfs with the near-IR spectro-polarimeter and velocimeter SPIRou at the Canada-France-Hawaii Telescope and with the optical velocimeter SOPHIE at the Haute-Provence Observatory in France. The aim of this paper is to present our results on Gl 410, a 0.55 M⊙ 480150 Myr old active M dwarf distant 12 pc.
Methods. We searched for planetary companions using radial velocities (RVs). We used the line-by-line (LBL) technique to measure the RVs with SPIRou and the template matching method with SOPHIE. Three different methods were employed, two based on principal component analysis (PCA), to clean the SPIRou RVs for systematics. We applied Gaussian processes (GP) modeling to correct the SOPHIE RVs for stellar activity. The and apodized sine periodogram analysis was used to search for planetary signals in the SPIRou data taking into account activity indicators. We analyzed TESS data and searched for planetary transits.
Results. We report the detection of a sin=8.41.3 M⊕ sub-Neptune planet at a period of 6.0200.004 days in circular orbit with SPIRou. The same signal, although with lower significance, was also retrieved in the SOPHIE RV data after correction for activity using a GP trained on SPIRou’s longitudinal magnetic field () measurements. The TESS data indicate that the planet is not transiting. Within the SPIRou wPCA RVs, we find tentative evidence for two additional planetary signals at 2.99 and 18.7 days.
Conclusions. Infrared RVs are a powerful method to detect extrasolar planets around active M dwarfs. Care should be taken, however, to correct or filter systematics generated by residuals of the telluric correction or small structures in the detector plane. The LBL technique combined with PCA offers a promising way to reach this objective. Further monitoring of Gl 410 is necessary.
@@
One of the most interesting open questions in current Astronomy is the census of extrasolar planets in the vicinity of the Solar System. The identification of the closest extrasolar planetary systems is a crucial milestone for the study of exoplanet atmospheres and biomarkers using future instrumentation combining high-dispersion spectroscopy with high contrast imaging (e.g., Kasper et al., 2021; Chauvin, 2024). As the most stars in the Galaxy, and thus in the Solar neighbourhood, are low-mass stars ( M⊙), in recent years, there has been growing interest in extending the planet search effort from solar-like stars to M dwarfs.
@@
Radial velocity (RV; Bonfils et al., 2013; Pinamonti et al., 2022; Mignon et al., 2025) and transit surveys (Dressing & Charbonneau, 2013, 2015; Mulders et al., 2015; Hsu et al., 2020; Ment & Charbonneau, 2023) have revealed that a large fraction of M dwarfs, if not all, host a planetary system, and these surveys have provided strong evidence supporting the hypothesis that the rocky planet ratio per star is indeed larger than one. Extrapolating the results from field stars to our stellar vicinity is extremely encouraging, as it suggests that it is highly likely that a large fraction of nearby M dwarfs host multi-planetary systems. Nearby stars are distributed all over the sky at random orientations. A small fraction of sources are expected to present planets in transit. Missions such as TESS, and in the future, PLATO will detect a significant fraction of those transiting planets. Nevertheless, to establish the planetary systems census of our closest stellar neighbors, techniques such as astrometry or RVs need to be used in addition to transits.
@@
Surveys for planetary systems around nearby bright M dwarfs with high-precision velocimeters have been carried so far mainly in the optical domain. Recently, the search using RVs has been extended to the near-infrared (IR) domain with the start of operations of new velocimeters such as HPF111Habitable Zone Planet Finder. (Mahadevan et al., 2014), CARMENES222Calar Alto high-Resolution search for M dwarfs with Exoearths with Near-infrared and optical Echelle Spectrographs. (Quirrenbach et al., 2018), SPIRou333SPectropolarimètre Infra-ROUge. (Donati et al., 2020) and NIRPS444Near Infra Red Planet Searcher. (Bouchy et al., 2017). High-precision IR RVs enable the exploration of lower mass stars, which are sometimes inaccessible for precise RV work in the optical due to their faintness. Furthermore, IR RVs have the potential of being less affected by stellar activity than optical RVs, especially for highly or moderately active stars (e.g., Carmona et al., 2023).
@@
With more than two thousand low-mass M dwarfs in the 25 pc vicinity, the effort to monitor all of these objects is colossal. Thanks to large programs with the near-IR spectra-polarimeter and velocimeter SPIRou at CFHT (the SPIRou Legacy Survey, the SPICE large program) and a series of PI SPIRou follow-up programs, our team has been monitoring a select sub-sample of around 60 nearby M dwarfs. At the time of writing, we have obtained more than 100 visits per source for around 45 of them.
@@
In this paper, we report the detection of a sub-Neptune mass planet in a 6-day period around the nearby (12 pc) 0.55 M⊙ 480 Myr old M dwarf Gl 410 with SPIRou, and the tentative detection of two additional planetary signals at 2.99 and 18.7 days. We complement our near-IR data set with quasi simultaneous optical observations obtained with the velocimeter SOPHIE555 Spectrographe pour l’Observation des Phénomènes des Intérieurs stellaires et des Exoplanètes (spectrograph for the observation of the phenomena of the stellar interiors and of the exoplanets).. The paper is organized as follows. We describe the physical properties of Gl 410 in Sect. 2. Then, in Sect. 3, we present the details of the optical and infrared RV measurements. In Sect. 4, we analyze our RV measurements and test the robustness of the planet detection. In Sect. 5, we search for the 6-day period signal detected with SPIRou in the SOPHIE data using Gaussian processes (GP) activity filtering techniques. We derive the orbital and physical properties of the detected 6.02 day planet using Markov chain Monte Carlo (MCMC) techniques in Sect. 6. In Sect. 7, we investigate the presence of other planetary signals within the SPIRou wPCA RVs. In Sect. 8, we present TESS observations of Gl 410 and investigate whether Gl 410b is transiting. Finally, we discuss our findings in Sect. 9, and in Sect. 10, we provide our conclusions.
@@
2 Gl410
We summarize the basic physical properties of Gl 410 (DS Leo, HD 95650) in Table 6. Gl 410 is a nearby star located at a distance of 11.9327 pc (Gaia DR3, Bailer-Jones et al., 2021; Gaia Collaboration et al., 2023). It has a spectral type M1.0V and a mass of 0.550.02 M⊙ (Cristofari et al., 2022). Gl 410 is known to exhibit cyclic photometric variability in the optical with periods of about 14 days, which is interpreted as being caused by rotational modulation of stellar spots (Fekel & Henry, 2000). The rotational period of Gl 410 has been tightly constrained using spectropolarimetry observations in the optical with ESPaDOnS, NARVAL, and HARPS (Donati et al., 2008; Hébrard et al., 2016), and in the near-IR with SPIRou (Fouqué et al., 2023; Donati et al., 2023). The latest estimation of is of 13.910.09 days (Donati et al., 2023). Gl 410 displays surface differential rotation with periods at the equator and pole of 13.370.86 and 14.961.25 days, respectively (Hébrard et al., 2016). Based on gyrochronologly Fouqué et al. (2023) derived an age of 0.890.1 Gyr for Gl 410. However, there is the possibility that Gl 410 is in fact younger. In Sect. A in the Appendix, we provide further details. We find that the empirical rotation-temperature sequences in clusters of known ages indicate that Gl 410 is 480150 Myr, thus placing it among the youngest stars in the solar neighborhood (Silva Aguirre et al., 2018; Chen et al., 2020) and suggesting that Gl 410 might be related to the Ursa Major (UMA) cluster which is approximately 400 Myr old.
Based on SPIRou observations, Cristofari et al. (2023) measured an average magnetic small scale field strength ¡¿ of 0.71 0.03 kG for Gl 410. Together with the M dwarfs GJ 1289 and GJ 1286, Gl 410 is one of the most magnetic stars among the M dwarfs studied in the ¡¿ surveys of Reiners et al. (2022) and Cristofari et al. (2023). The magnetic activity of Gl 410 is further indicated by the presence of emission in the optical spectra. Gl 410 is part of the top one-third of the sample in terms of emission strength among the 331 M stars surveyed by Schöfer et al. (2019).
Zeeman Doppler imaging (ZDI) investigations by Donati et al. (2008), Hébrard et al. (2016), and Bellotti et al. (2024) have shown that the large-scale magnetic field of Gl 410 rapidly evolves both in strength and in topology. Between 2007 and 2008, the field had a predominantly toroidal geometry with an average strength of 100 G (Donati et al., 2008). In 2014, the field remained mostly toroidal with axisymmetric geometry, but the average field strength decreased to 60 G (Hébrard et al., 2016). In 2020b to 2021a the magnetic energy became almost equally distributed between the poloidal and the toroidal components and the average field strength decreased to 44 G (Bellotti et al., 2024). In the latest measurements from 2021b to 2022a, the poloidal component became much stronger than the toroidal component with 73% of the magnetic energy, and the average field strength further decreased to 18 G (Bellotti et al., 2024).
@@ 3.1 SOPHIE
Gl 410 was observed during two periods 9 years apart with the high-resolution optical velocimeter SOPHIE (Perruchot et al., 2008; Bouchy et al., 2013) at the 1.93m Telescope of the Haute-Provence Observatory in the South of France. The first period with 45 visits was observed from March 14, 2010, to June 14, 2012. A second observation period was planned to coincide with the SPIRou observations, with 58 visits performed from January 26, 2021, to March 04, 2023. Spectra were reduced using SOPHIE’s data reduction system (Bouchy et al., 2009; Heidari, 2022). The RV measurements were obtained using a template-matching technique implemented in the code NAIRA (Astudillo-Defru et al., 2015), which has been adapted for SOPHIE (Hobson et al., 2018). The SOPHIE RV measurements are given in Table 8 (the full table is available at the CDS). In Table 4, we provide a summary of the properties of the RV time series. Figure 1 displays the median subtracted SOPHIE RV data.
@@ 3.2 SPIRou
Gl 410 was observed with the near-IR velocimeter and polarimeter SPIRou (Donati et al., 2020), mounted at the Canada-France-Hawaii Telescope atop Mauna-Kea during 156 visits from April 2018 to July 2023. Data were obtained in the context of the SPIRou’s commissioning in 2018, the SPIRou Legacy Survey from 2019 to 2022A, and a series of PI follow-program programs (PI. Carmona, PI. Artigau) in 2022B, 2023A and 2023B. Nightly SPIRou observations typically consist of a Stokes V sequence of four exposures, each taken with different positions of the Fresnel rhombs inside the polarimeter. The combination of the channel A and channel B of the four sub-exposures enables the recovery of the Stokes V (and Stokes I) spectra for magnetic field analysis. For each science observation, a simultaneous Fabry-Pérot exposure was obtained in the calibration channel C. In the context of the RV analysis, we used the combined extraction of the science channels A and B and each sub-exposure of the Stokes V sequence was treated independently.
For the RV analysis, we used the data reduced using the APERO (Cook et al., 2022) pipeline version v07275 installed at the SPIRou Data Centre hosted at the Laboratoire d’Astrophysique de Marseille. We note that APERO extracts the combined science channel A and channel B spectra, performs the wavelength calibration using nightly observations of the Uranium-Neon lamp and Fabry-Pérot exposures, and conducts the telluric correction. The telluric correction uses a combination of an atmospheric absorption model generated by TAPAS (Bertaux et al., 2014), an empirical template of the star deduced from observations at different barycentric earth radial velocity (BERV) and a library of nightly observations of telluric standard stars taken since the commissioning of the instrument in 2018. Further details on the telluric correction are given in Artigau et al. (2014), Cook et al. (2022) and Artigau et al. (in prep.). Individual RV measurements were obtained using the LBL technique (Artigau et al., 2022), and they were corrected by the BERV and the drift of the Fabry-Pérot at each exposure. A further correction for the instrumental RV zero point was done using a GP model based on observations of a sample of RV standard stars observed by SPIRou (Cadieux et al., 2022). The SPIRou LBL RV measurements are given in Table 9 (full table available at the CDS).
For analysis of the stellar activity, we used the simultaneous longitudinal magnetic () field measurements obtained with the Libre-ESpRIT pipeline777The nominal ESPaDOnS data reduction pipeline (Donati et al., 1997) adapted for the SPIRou data (Donati et al., 2020). previously published in Donati et al. (2023).
| DATE | RJD | RV | RV | ||
| [yyyy-mm-dd] | [days] | [km/s] | [km/s] | [m/s] | [m/s] |
| 2010-03-14 | 55269.43680 | -13.8604 | 0.0017 | -21.9 | 1.7 |
| 2010-03-15 | 55270.60680 | -13.8701 | 0.0026 | -31.6 | 2.6 |
| 2010-03-16 | 55271.58010 | -13.8473 | 0.0027 | -8.8 | 2.7 |
| LBL | Wapiti | |||||||||
| DATE | RJD | RV | RV | RV | RV | |||||
@@ 3.3 TESS
The Transit Exoplanet Survey (TESS; Ricker et al., 2014) observed Gl 410 (TIC 97488127) with a two-minute cadence during Sector 22, from February 18 to March 18, 2020. We obtained the Presearch Data Conditioning (PDC) flux time series (Smith et al., 2012; Stumpe et al., 2012, 2014) processed by the TESS Science Processing Operations Center (Jenkins et al., 2016) from the Mikulski Archive for Space Telescopes (MAST)101010mast.stsci.edu.